# A Smooth Transition from Giant Planets to Brown Dwarfs from the Radial Occurrence Distribution

### v1

@@ 5.2 The Brown Dwarf Desert
We showed in Section 4.2 that occurrence decreases with increasing mass, reaching a minimum for brown dwarfs (13–80 $M_{\mathrm{Jup}}$). This interval extends to a mass ratio ($q=M_{c}/M_{\text{pri}}$) of approximately $q=0.1$ for the Sun-like stars that comprise most of the CLS. Although we were unable to measure occurrence rates for higher-mass (stellar) companions, studies of stellar multiplicity over wide ranges in separation and primary type have found that the occurrence rate increases with increasing mass ratio for low $q$ ($\sim$0.1–0.3; Reggiani2013; Moe2017). Assuming this pattern holds for solar hosts between 1–10 AU, the 13–80 $M_{\mathrm{Jup}}$ interval represents a valley in the companion occurrence distribution, meaning that the brown dwarf desert found by GretherLineweaver2006 for separations $\lesssim$3 AU reaches, and perhaps exceeds, 10 AU.
@@
The extension of the brown dwarf desert to at least 10 AU has ramifications for theories of brown dwarf dynamical evolution. In addition to a brown dwarf desert out to 5-year orbital periods ($\sim$3 AU), GretherLineweaver2006 reported a lack of such a desert among free-floating BDs in the Orion, Pleiades, and M35 stellar clusters. In other words, brown dwarfs often form in isolation, but rarely as companions to Sun-like stars. They argued that this distinction disfavors a minimum in the efficiency of gravitational collapse or disk fragmentation (IdaLin2004-metallicity) as the source of the BD desert, and instead supports post-formation migration as the mechanism preventing the settlement of BDs at close separations. Our findings are consistent with this picture, but require that such migrations mechanisms be capable of explaining the BD desert out to 10 AU.
5.3 Comparison to Direct Imaging Surveys
@@
The CLS’s achievement of sensitivity to companions near 10 AU is an important bridge between the short- and long-period regimes so far probed by RV and imaging surveys, respectively. It has allowed for comparison between occurrence rates calculated by these two methods (Fulton2021), under the assumption that giant planet occurrence remains constant between young ($\sim$10 Myr) and intermediate (Gyr) stellar ages.
@@
In this work, we addressed a remaining barrier to direct comparison, which is the discrepancy between the minimum masses measured by RVs alone and the true masses estimated from direct imaging through companion formation and luminosity models (Baraffe03; Nielsen2019). We measured the occurrence of companions between 0.03–30 AU and 0.8–80 $M_{\mathrm{Jup}}$, both as functions of mass and separation together (Section 4.1) and of each parameter separately (Sections 4.2 and 4.3).
@@
In Section 4.2, we found an occurrence rate of $0.9^{+0.5}_{-0.4}\%$ for brown dwarfs ($M_{c}$=13–80 $M_{\mathrm{Jup}}$) between 1–10 AU. Assuming the occurrence rate is log-constant at wide separations, this value is in close agreement with that of Nielsen2019, who estimated a BD occurrence rate of $0.8^{+0.8}_{-0.5}\%$ between 10–100 AU from observations made with the Gemini Planet Imager (Macintosh2014). If we extrapolate our BD occurrence rate between 10–30 AU ($1.7^{+1.0}_{-0.7}\%$) to 10–100 AU, assuming a log-uniform rate, we arrive at $3.4^{+2.0}_{-1.4}$%, larger than the estimate of Nielsen2019. We note a number of caveats in our comparison: the large fractional uncertainties in both measurements, the difference in primary star masses and ages in the two samples, and the uncertainty in the period distribution of brown dwarfs between 10–100 AU.
5.4 Biases
@@
Our study is subject to multiple biases, which we list here. First, we incorporated HGCA data into our fits only for systems exhibiting significant astrometric accelerations. Because larger companions produce more significant accelerations, our strategy allowed us to identify massive inclined companions, but was unable to precisely constrain the inclinations of their low-mass counterparts. This is analogous to the RV bias against low-mass, wide-separation objects, which we accounted for using a system-by-system sensitivity correction. A detailed accounting of the HGCA’s sensitivity would improve our result by accounting for the small inclined objects that the HGCA missed.
@@
Another limitation in our analysis is the exclusion of RV trends. Long-period companions produce non-periodic signals that are consistent with a wide range of companion models, often including planets, brown dwarfs, and stars. Rosenthal2021 reported trends in 89 systems, nearly as many as the number of systems hosting companions with full orbits. Although it is possible to model such signals probabilistically (e.g. VanZandt2024; VanZandt2025), we opted not to include them in our analysis due to their comparatively minor contribution to occurrence statistics in our parameter space of interest. A future study incorporating trends would strengthen occurrence constraints, particularly at wide separations.
6 Conclusion
@@
We re-fit the orbits of 192 substellar and 3 stellar companions from the California Legacy Survey to measure occurrence at wide separations ($\lesssim$30 AU). We combined the multi-decade RV data sets of the California Legacy Survey with absolute astrometry from the HipparcosHipparcos-GaiaGaia Catalog of Accelerations to fit three-dimensional orbits where possible. We identified eight of the 18 apparently-substellar companions ($M_{c}\sin i$ = 13–80 $M_{\mathrm{Jup}}$) in the CLS catalog as inclined M dwarfs.
@@
  • The lack of a sharp transition between either the shape of the semi-major axis distributions of objects between 0.8–80 $M_{\mathrm{Jup}}$ (Section 4.1), or the marginalized occurrence rate between 1–10 AU (Section 4.2), disfavors a “cut-off mass” below which planets are exclusively formed by core accretion and above which brown dwarfs are exclusively formed by gravitational instability. Instead, the smooth variation in both distributions suggests that the two mechanisms produce companions in overlapping mass ranges.
  • Companion occurrence between 1–10 AU is a monotonically decreasing function of mass, reaching a minimum for brown dwarfs (13–80 $M_{\mathrm{Jup}}$), consistent with previous findings of a ’brown dwarf desert’ at shorter separations. Whatever mechanism(s) stifle the formation/migration of brown dwarfs at low separations are relevant at wider separations as well.
  • The radial occurrence distribution of all companions between 0.8–80 $M_{\mathrm{Jup}}$ exhibits a significant preference for a step near 1 AU, consistent with Fulton2021’s finding of an enhancement in the occurrence of companions with $M\sin i$= 30–6000 $M_{\oplus}$ beyond 1 AU. Determining precisely which mass sub-intervals drive this feature will help determine the dividing line between companions generated by core accretion versus gravitational/disk instability. The possibility of a drop-off in occurrence beyond 10 AU persists as well, though the large uncertainties at these wide separations prevent a firm conclusion.

### v2

@@ V.2 The Brown Dwarf Desert
We showed in Section IV.2 that occurrence decreases with increasing mass, reaching a minimum for brown dwarfs (13–80 $M_{\mathrm{Jup}}$). This interval extends to a mass ratio ($q=M_{c}/M_{\text{pri}}$) of approximately $q=0.1$ for the Sun-like stars that comprise most of the CLS. Although we were unable to measure occurrence rates for higher-mass (stellar) companions, studies of stellar multiplicity over wide ranges in separation and primary type have found that the occurrence rate increases with increasing mass ratio for low $q$ ($\sim$0.1–0.3; Reggiani and Meyer 2013; Moe and Di Stefano 2017). Assuming this pattern holds for solar hosts between 1–10 AU, the 13–80 $M_{\mathrm{Jup}}$ interval represents a valley in the companion occurrence distribution, meaning that the brown dwarf desert found by Grether and Lineweaver (2006) for separations $\lesssim$3 AU reaches, and perhaps exceeds, 10 AU. This feature was observed recently by An et al. (2025), who acknowledged that their sample was not designed for occurrence rate calculations. Its presence in both samples provides strong evidence that the brown dwarf desert reaches 10 AU.
@@
The extension of the brown dwarf desert to at least 10 AU has ramifications for theories of brown dwarf dynamical evolution. In addition to a brown dwarf desert out to 5-year orbital periods ($\sim$3 AU), Grether and Lineweaver (2006) reported a lack of such a desert among free-floating BDs in the Orion, Pleiades, and M35 stellar clusters. In other words, brown dwarfs often form in isolation, but rarely as companions to Sun-like stars. They argued that this distinction disfavors a minimum in the efficiency of gravitational collapse or disk fragmentation (Ida and Lin, 2004) as the source of the BD desert, and instead supports post-formation migration as the mechanism preventing the settlement of BDs at close separations. Our findings are consistent with this picture, but require that such migrations mechanisms be capable of explaining the BD desert out to 10 AU.
V.3 Comparison to Direct Imaging Surveys
@@
The CLS’s achievement of sensitivity to companions near 10 AU is an important bridge between the short- and long-period regimes so far probed by RV and imaging surveys, respectively. It has allowed for comparison between occurrence rates calculated by these two methods (Fulton et al., 2021), under the assumption that giant planet occurrence remains constant between young ($\sim$10 Myr) and intermediate (Gyr) stellar ages.
@@
In this work, we addressed a remaining barrier to direct comparison, which is the discrepancy between the minimum masses measured by RVs alone and the true masses estimated from direct imaging through companion formation and luminosity models (Baraffe et al., 2003; Nielsen et al., 2019). We measured the occurrence of companions between 0.03–30 AU and 0.8–80 $M_{\mathrm{Jup}}$, both as functions of mass and separation together (Section IV.1) and of each parameter separately (Sections IV.2 and IV.3).
@@
In Section IV.2, we found an occurrence rate of $1.1^{+0.5}_{-0.4}\%$ for brown dwarfs ($M_{c}$=13–80 $M_{\mathrm{Jup}}$) between 1–10 AU. Assuming the occurrence rate is log-constant at wide separations, this value is in close agreement with that of Nielsen et al. (2019), who estimated a BD occurrence rate of $0.8^{+0.8}_{-0.5}\%$ between 10–100 AU from observations made with the Gemini Planet Imager (Macintosh et al., 2014). If we extrapolate our BD occurrence rate between 10–30 AU ($1.6^{+0.9}_{-0.6}\%$) to 10–100 AU, assuming a log-uniform rate, we arrive at $3.2^{+1.3}_{-0.8}$%, larger than the estimate of Nielsen et al. (2019). We note a number of caveats in our comparison: the large fractional uncertainties in both measurements, the difference in primary star masses and ages in the two samples, and the uncertainty in the period distribution of brown dwarfs between 10–100 AU.
V.4 Distribution of Orvara Inclination Draws
Following An et al. (2025), we validated our Orvara fits by evaluating the distribution of posterior orbital inclination draws among all companion fits. Assuming an unbiased sample, we expect the distribution of sampled inclinations to match the geometric inclination prior, that is, the distribution that arises from an ensemble of orbits randomly oriented in three dimensions. A deviation from this prior could indicate systematic biases in our sample selection or obit fitting procedures. After combining all MCMC chains across all fitted companions, we find that the inclination distribution fits a sinusoid well. Additionally, we examined the inclination distribution of the companions in the 53 systems in our sample exhibiting a significant HGCA acceleration ($\Delta\mu/\sigma_{\Delta\mu}\geq 3$). We found that this distribution was likewise sinusoidal, though with greater deviation as compared to the full sample. This deviation is expected, given that filtering on astrometric acceleration significance introduces a preference for massive companions on inclined orbits. Figure 5 depicts the inclination distributions of our full and astrometrically accelerating samples.
Figure 5: Left: Posterior distribution of inclinations from all Orvara chains in our sample, including all companions in multi-companion systems. The red line shows a sinusoidal distribution, which a large sample of randomly-oriented companion inclinations is expected to follow. We scaled this distribution so that its integral equals that of the posterior. Right: The same as the right panel, but for only the 53 systems in which we calculated a significant ($\geq 3\sigma$) astrometric acceleration. As expected, we find greater deviation from the theoretical distribution among this sample, which is biased to include massive companions on inclined orbits. Nevertheless, the roughly sinusoidal distribution of even these inclinations suggests that our fitting procedure did not introduce heavy bias to our analysis.
V.5 Biases
@@
Our study is subject to multiple biases, which we list here. First, as noted in Section V.4, we incorporated HGCA data into our fits only for systems exhibiting significant astrometric accelerations. Because larger companions produce more significant accelerations, our strategy allowed us to identify massive inclined companions, but was unable to precisely constrain the inclinations of their low-mass counterparts. This is analogous to the RV bias against low-mass, wide-separation objects, which we accounted for using a system-by-system sensitivity correction. Additionally, as noted by An et al. (2025), the RV bias toward edge-on orbits and the astrometric bias toward face-on orbits implies that a combination of these two techniques can mitigate the overall inclination bias. Nevertheless, a detailed accounting of the HGCA’s sensitivity would improve our result by accounting for the small inclined objects that the HGCA missed.
@@
Another limitation in our analysis is the exclusion of RV trends. Long-period companions produce non-periodic signals that are consistent with a wide range of companion models, often including planets, brown dwarfs, and stars. Rosenthal et al. (2021) reported trends in 89 systems, nearly as many as the number of systems hosting companions with full orbits. Although it is possible to model such signals probabilistically (e.g. Van Zandt and Petigura 2024a; Van Zandt et al. 2025), we opted not to include them in our analysis due to their comparatively minor contribution to occurrence statistics in our parameter space of interest. A future study incorporating trends would strengthen occurrence constraints, particularly at wide separations.
VI Conclusion
@@
We re-fit the orbits of 191 substellar and 3 stellar companions from the California Legacy Survey to measure occurrence at wide separations ($\lesssim$30 AU). We combined the multi-decade RV data sets of the California Legacy Survey with absolute astrometry from the HipparcosHipparcos-GaiaGaia Catalog of Accelerations to fit three-dimensional orbits where possible. We identified seven of the 18 apparently-substellar companions ($M_{c}\sin i$ = 13–80 $M_{\mathrm{Jup}}$) in the CLS catalog as inclined M dwarfs.
@@
  • The lack of a sharp transition between either the shape of the semi-major axis distributions of objects between 0.8–80 $M_{\mathrm{Jup}}$ (Section IV.1), or the marginalized occurrence rate between 1–10 AU (Section IV.2), disfavors a “cut-off mass” below which planets are exclusively formed by core accretion and above which brown dwarfs are exclusively formed by gravitational instability. Instead, the smooth variation in both distributions suggests that the two mechanisms produce companions in overlapping mass ranges.
  • Companion occurrence between 1–10 AU is a monotonically decreasing function of mass, reaching a minimum for brown dwarfs (13–80 $M_{\mathrm{Jup}}$), consistent with previous findings of a “brown dwarf desert” at shorter separations. Whatever mechanism(s) stifle the formation/migration of brown dwarfs at low separations are relevant at wider separations as well.
  • The radial occurrence distribution of all companions between 0.8–80 $M_{\mathrm{Jup}}$ exhibits a significant preference for a step near 1 AU, consistent with Fulton et al. (2021)’s finding of an enhancement in the occurrence of companions with $M\sin i$= 30–6000 $M_{\oplus}$ beyond 1 AU. Determining precisely which mass sub-intervals drive this feature will help determine the dividing line between companions generated by core accretion versus gravitational/disk instability. The possibility of a drop-off in occurrence beyond 10 AU persists as well, though the large uncertainties at these wide separations prevent a definitive conclusion.